# Don’t Let It Hallucinate: Premise Verification via Retrieval-Augmented Logical Reasoning

### v1

@@ 4.2.3 Hallucination Mitigation Baselines
2) Prompt: We encourage the LLM to assess potential false premises before generating a response by appending the following prompt to the original query:
This question may contain a fasle premise. [query]
3) Majority Vote: We prompt the LLM three times with the same prompt and select the most frequent response as the final answer. This method improves reliability by reducing the impact of any single erroneous or hallucinated response from LLM.
4) Perplexity AI [^3] https://www.perplexity.ai: Utilizes a search engine to retrieve and incorporate real-time information from the web, enabling it to provide answers based on the latest available web data.
For Direct Ask and Majority Vote, we report the performances of the following LLMs: GPT-3.5-turbo (OpenAI, 2023), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), and Qwen-1.5-7b-chat (Bai et al., 2023).
Models Direct Ask Our Method Prompt Majority Vote Perplexity AI
GPT-3.5 93.3 94.3$\uparrow$ 93.3 92.4 92.4
Mistral-7B 87.6 89.5$\uparrow$ 86.7 87.6 92.4
Qwen1.5 89.5 91.4$\uparrow$ 90.5 90.5 92.4
Table 2: Comparison of accuracy of different hallucination mitigation methods.
4.3 Metrics
In the false premise detection task, we look at several metrics: the true positive rate (TPR), the true negative rate (TNR), the false positive rate (FPR), the false negative rate (FNR), the F1 score, and accuracy of the model successfully identifying questions containing false premises or not. Here, a positive instance refers to a question that contains a false premise, meaning a higher TPR indicates better detection of false premises. For the hallucination mitigation task, we focus on question-answering accuracy. We calculate accuracy by string matching the responses of LLMs: for TPQs, answering “Yes” is considered correct; for FPQs, answering “No” is considered correct.
@@ 5 Discussion
We show the result of the false premise detection task in Table 1 and the hallucination mitigation result in Table 2.
@@
Using logical forms helps better identify false premises in the questions. As shown in Table 1, for all three retrievers, explicitly incorporating logical forms into both retrieval and false premise detection stages significantly improves the identification of false premises. Sole reliance on original queries, even though potentially yielding high accuracy, tends to neglect accurate false premise identification, underscoring the importance of utilizing structured logical forms for tasks prioritizing precise false premise detection.
@@
Among different types of retrievers when using logical forms in both the retrieval and false premise detection stages, the G-retriever method achieves the highest TPR at 75.56%, demonstrating a strong capability in accurately identifying questions containing false premises. Notably, this method also achieves the highest F1 score (84.47%), indicating an optimal balance between precision and recall. Although the ToG method exhibits the highest TNR of 93.33%, it significantly underperforms in TPR and overall F1 score (16.16%), suggesting limited effectiveness in correctly identifying false premises.
@@
Notably, when original queries are used in either retrieval, false premise detection, or both stages, despite achieving high accuracy (91.11%), G-retriever shows a markedly lower TPR (37.78%) compared to the first configuration. This suggests that relying on original queries alone, or in combination with logical forms in only one stage for detection, can achieve high accuracy due to correctly identifying negatives, it is less effective at capturing false premises, which is the primary focus of our task.
@@
Explicitly detecting and informing LLMs false premise mitigates hallucination, as demonstrated in Table 2. Our proposed method, which directly communicates the presence of false premises to the models, achieves the highest accuracy: 94.3% with GPT-3.5, 91.4% with Qwen1.5, and 87.6 % with Mistral-7B. This performance surpasses alternative approaches such as Direct Ask, Prompt, Majority Vote, and Perplexity AI, which yield lower accuracy.
@@
Majority Vote does not perform well, likely due to hallucination snowballing, where repeated querying amplifies errors rather than correcting them. Additionally, while the Prompt method warns the model about potential false premises, it does not specifically tell the LLM which one contains false premises, negatively impacts performance on questions with valid premises, causes unnecessary cautiousness and reduces the model’s ability to provide direct and confident answers. Besides, Perplexity AI does not perform as well potentially because the query format does not align well with web data, leading to suboptimal retrieval of relevant information for certain types of questions. These findings emphasize the importance of tailoring hallucination mitigation strategies to both the model’s reasoning process and the nature of the queries it encounters.
@@
Figure 4: GPT-3.5-turbo: False premise detection accuracy across single-hop and multi-hop queries. Using logical form-based RAG mainly helps detect false premises in multi-hop questions.
@@
Our approach mostly improves false premise detection performance on multi-hop questions, according to Figure 4. The incorporation of logical form-based RAG leads to notable performance gains compared to direct claim evaluation. Specifically, while single-hop questions see moderate improvement, multi-hop questions benefit the most, with false premise detection performance increasing from 48.3% in the direct claim setting to 66.7% with RAG and further to 73.3% when using the G-retriever. These results suggest that leveraging retrieval mechanisms enhances reasoning over multiple pieces of evidence, reinforcing the importance of retrieval-augmented methods for complex question-answering tasks. We present a case study to illustrate how our method improves performance on multi-hop questions in Appendix A.2.
@@ 6 Conclusion
In this work, we propose a retrieval-augmented logical reasoning framework to detect false premises and then mitigate hallucinations in LLMs. Our method explicitly detects and signals false premises, overcoming key limitations of current approaches that rely on model parameters or post-hoc corrections. By incorporating explicit false premise detection, we effectively mitigate hallucinations without requiring output generation or direct access to model logits. Our results demonstrate that logical forms significantly improve the identification of false premises, particularly in multi-hop questions where reasoning over multiple steps is required. The proposed approach enhances LLM robustness by providing a structured mechanism to detect and handle misleading inputs before they influence downstream responses. This reinforces the importance of structured reasoning techniques in improving model reliability and factual consistency.
Limitations and Future Work. Our work has some limitations. Firstly, the performance improvement varies across different retrievers and LLMs, and we do not cover all the situations. As noted in (Zhu et al., 2024), some models such as Llama and Baichuan2 series may have an inherent bias that causes them to consistently favor one type of answer when answering Yes-No questions. Secondly, our method requires a prior knowledge graph, which may not always be feasible and may be difficult to update in line with developments in the real world. Lastly, we look at the factuality hallucination caused by false premise in discriminative tasks, which does not encompass all possible tasks. Future work can explore broader applications of logical form-based detection across diverse reasoning tasks and investigate integration strategies that further align LLM behavior with verifiable logical structures.
Broader Impact Statement. Our work addresses hallucinations in LLMs by detecting and mitigating false premises prior to response generation. This approach enhances reliability, reduces misinformation, and promotes responsible AI use. Ethical considerations, including fairness and bias, should be carefully addressed in future applications.
@@ A.1 Dataset Details
A.2 Case Study

### v2

@@ 4.2.3 Hallucination Mitigation Methods
2) Prompt: We encourage the LLM to assess potential false premises before generating a response by appending the following prompt to the original query: This question may contain a fasle premise. [query]
3) Majority Vote (MajVote): We prompt the LLM three times with the same prompt and select the most frequent response as the final answer. This method improves reliability by reducing the impact of any single erroneous or hallucinated response. from LLM.
4) Perplexity AI [^3] https://www.perplexity.ai: Utilizes a search engine to retrieve and incorporate real-time information from the web, enabling it to provide answers based on the latest available web data. We use the version powered by GPT-4-Omni.
5) Direct RAG: Retrieves relevant entities from the knowledge graph and provides them as context to the LLM alongside the original query. This approach augments the model’s internal knowledge directly with external information to improve answer accuracy and grounding.
We report the performances of the following LLMs: GPT-4o-mini (OpenAI, 2024), GPT-3.5-turbo (OpenAI, 2023), Llama-3.1-8B-Instruct (et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Qwen2.5-7B-Instruct (Qwen et al., 2025), and Qwen-1.5-7b-chat (Bai et al., 2023). To better understand where the performance gains come from, we provide a per-model breakdown of error corrections and premise-specific changes (false- vs. true-premise queries) in Appendix §A.4.1.
Models DirectAsk Prompt MajVote DirectRAG Ours
GPT-4o-mini 83.8 92.4 86.7 90.5 92.4
GPT-3.5 93.3 93.3 92.4 90.5 94.3
LLama-3.1 86.7 86.7 89.5 88.6 89.5
Mistral-7B 87.6 86.7 87.6 71.4 89.5
Qwen2.5 92.4 86.7 92.4 92.4 95.2
Qwen1.5 89.5 90.5 90.5 82.9 91.4
Perplexity AI 91.4
Table 6: Comparison of accuracy (%) of different hallucination mitigation methods.
Metrics. We evaluate question-answering accuracy on the hallucination mitigation task. Accuracy is calculated by string matching the responses of LLMs: for TPQs, answering “Yes” is considered correct; for FPQs, answering “No” is considered correct.
@@ 5 Discussion
We show the result of the false premise detection task in Tab.2, 3, and 4 for the KG-FPQ, CREAK and FEVER dataset, respectively. The SAC3 baseline result is shown in Appendix § A.4.5. Tab. 6 presents the hallucination mitigation result.
@@
Using logical forms helps better identify false premises in the questions. As shown in Tab. 2, for all three retrievers, explicitly incorporating logical forms into both retrieval and false premise detection stages significantly improves the identification of false premises. Sole reliance on original queries, even though potentially yielding high accuracy, tends to neglect accurate false premise identification, underscoring the importance of utilizing structured logical forms for tasks prioritizing precise false premise detection.
@@
For the KG-FPQ dataset, among different types of retrievers, when using logical forms in both the retrieval and false premise detection stages, the G-retriever method achieves the highest TPR at 94.44%, demonstrating a strong capability in accurately identifying questions containing false premises. Notably, this method also achieves the highest F1 score (97.12%), indicating an optimal balance between precision and recall. Although the ToG method exhibits the highest TNR of 93.33%, it significantly underperforms in TPR and overall F1 score (16.16%), suggesting limited effectiveness in correctly identifying false premises.
@@
Notably, when original queries are used in either retrieval, false premise detection, or both stages, despite achieving reasonable accuracy (73.33% and 79.69%), with RAG method shows significantly lower TPR (33.33% and 37.78%) compared to the first configuration. This suggests that relying on original queries alone, or in combination with logical forms in only one stage for detection, can achieve high accuracy due to correctly identifying negatives, it is less effective at capturing false premises, which is the primary focus of our task.
Similarly, for the CREAK dataset, according to Tab. 3, using logical forms in both retrieval and detection stages consistently boosts performance versus operating on the original query. The gains are most pronounced for G-retriever: F1 score increases from 71.1% to 90.4% and accuracy from 74.5% to 90.5%, driven by a large increase in TPR (62.3% to 88.4%). GraphRAG/ToG already performs strongly with original queries, but still benefits from using logical forms, reaching the best overall scores (F1 92.1% and Acc 92.0%) with higher TPR (89.9% to 92.8%) while keeping FPR modest. The mixed configuration for retrieval mainly helps G-retriever but is less reliable than using logical forms in both stages, underscoring that structured queries plus explicit false premise detection are jointly necessary. Logical forms improves evidence targeting and makes contradictions salient to LLMs. Overall, the results indicate that logic-aware framework is crucial for converting strong retrieval into consistent end-to-end factuality.
According to Tab. 4, the proposed design also yields consistent performance gains on natural language knowledge base, indicating that incorporating logical forms in both the premise detection and LLM response stages remains effective when the knowledge source is less structured. This suggests that the benefits of logical-form guided reasoning stem from improved premise understanding and response control, rather than reliance on a specific knowledge representation.
@@
Explicitly detecting and informing LLMs false premise mitigates hallucination, as demonstrated in Tab. 6. Our proposed method, which directly communicates the presence of false premises to the models, achieves the highest accuracy: 92.4% with GPT-4o-mini, 94.3% with GPT-3.5, 95.2% with Qwen2.5, and 91.4% with Qwen-1.5. This performance surpasses alternative approaches such as Direct Ask, Prompt, Majority Vote, DirectRAG, and Perplexity AI.
@@
Majority Vote does not perform well, likely due to hallucination snowballing, where repeated querying amplifies errors rather than correcting them. Additionally, while the Prompt method warns the model about potential false premises, it does not specifically tell the LLM which one contains false premises, negatively impacts performance on questions with valid premises, causes unnecessary cautiousness and reduces the model’s ability to provide direct and confident answers. Besides, Perplexity AI does not perform as well potentially because the query format does not align well with graph data, leading to suboptimal retrieval of relevant information for certain types of questions. These findings emphasize the importance of tailoring hallucination mitigation strategies to both the model’s reasoning process and the nature of the queries.
Direct RAG retrieves and feeds raw evidence to LLMs without explicitly structuring the reasoning process or highlighting inconsistencies between the query and retrieved facts. As a result, the model may surface relevant but semantically unaligned information, leading to shallow retrieval-based responses rather than true logical verification. In contrast, logical-form RAG + explicit false premise signaling forms a structured reasoning process: it decomposes the claim into logical predicates and explicitly indicates when a premise conflicts with retrieved evidence. This guides LLMs to perform fact-level reasoning and contradiction handling, reducing overreliance on surface overlap and improving factual precision and interpretability.
Beyond hallucination mitigation, the proposed premise verification mechanism can be extended to sensitive or controversial topics, where unverified premises may amplify misinformation or harmful narratives. By explicitly detecting unsupported assumptions prior to response generation, the method offers a principled way to prevent models from uncritically engaging with inaccurate or inflammatory premises, enabling safer and more grounded interactions in high-risk domains. This suggests a broader role for retrieval-augmented logical reasoning as a lightweight safeguard for responsible deployment, especially in scenarios where factual grounding is essential before engaging in downstream reasoning or dialogue.
@@
Figure 4: GPT-4o-mini and G-retriver: False premise detection accuracy across single-hop and multi-hop queries on the KG-FPQ dataset. Using logical form-based RAG mainly helps detect false premises in multi-hop questions.
@@
Our approach mostly improves false premise detection performance on multi-hop questions, according to Fig. 4. The incorporation of logical form-based RAG leads to notable performance gains compared to direct claim evaluation. Specifically, while single-hop questions see moderate improvement, multi-hop questions benefit more, with false premise detection performance increasing from 68.7% in the direct claim setting to 79.3% with RAG and further to 95.2% when using the G-retriever. These results suggest that leveraging retrieval mechanisms enhances reasoning over multiple pieces of evidence, reinforcing the importance of retrieval-augmented methods for complex question-answering tasks. We present a case study to illustrate how our method improves performance on multi-hop questions in Appendix § A.4.6.
5.1 Computational Cost Analysis
In Tab. 5, we compare our method with the post-hoc Contrastive Decoding (Shi et al., 2023) approach in terms of computational efficiency and model compatibility (accuracy result based on Llama-3.1-8B). Our method reduces running time, uses fewer tokens by leveraging logical forms, and supports both model-agnostic and black-box settings. The proposed method introduces only a modest computational overhead, adding $\mathcal{O}(k+s)$ complexity to the retriever, where $k$ is the number of retrieved candidates and $s$ is the feature size used for logical-form reasoning and premise detection. In practice, this addition remains lightweight since most of the cost lies in the embedding model inference, which requires approximately 0.335 TFLOPs per example, while the remaining steps—retrieval, logical form conversion, and false-premise detection—are implemented as efficient API calls. In contrast, post-hoc methods rely on fine-tuning and lack general applicability across different model architectures. We also include performance comparison of Contrastive Decoding with other LLMs in Appendix § A.4.3.
5.1.1 Significance Test
To ensure that the reported performance differences are statistically meaningful, we conduct paired t-tests across model configurations. The results, summarized in Table 8 and Table 8 , confirm that the observed improvements are statistically significant (p < 0.05) in most comparisons. Table 8 evaluates overall differences among performances of Direct Claim, with RAG, and G-Retriever on the KG-FPQ dataset. Table 8 focuses on intra-method variations, contrasting setups that use logical forms in both stages (GG), only one stage (GO), or none (OO). The results confirm that applying logical forms consistently across both retrieval and detection stages yields significantly higher performance, reinforcing the benefit of structured reasoning input.
Comparison p-value
G-Retriever – wRAG 0.04
Direct Claim – G-Retriever 0.001
wRAG – G-Retriever 0.02
Table 7: Statistical significance comparison across retrieval methods.
Comparison p-value
G-Retriever (GG) – G-Retriever (GO) <0.001<0.001
G-Retriever (GG) – G-Retriever (OO) <0.001<0.001
wRAG (GG) – wRAG (GO) 0.01
Table 8: Significance comparison of configurations within retrieval methods.
@@ 6 Conclusion
We propose a retrieval-augmented logical reasoning framework that detects false premises to mitigate LLM hallucinations. Our method explicitly detects and signals false premises, overcoming key limitations of current approaches that rely on model parameters or post-hoc corrections. By incorporating upfront false premise detection, we prevent hallucinations without requiring output generation or model logit access. Results show logical forms significantly improve false premise identification, especially for multi-hop reasoning questions. Our approach enhances LLM robustness by providing a structured mechanism to detect and handle misleading inputs before they influence downstream responses. This reinforces the importance of structured reasoning techniques in improving model reliability.
@@ A.1 KG-FPQ Dataset Details
A.2 CREAK Dataset Details
The CREAK dataset Onoe et al. (2021) is designed to test whether language models can combine factual knowledge about specific entities with commonsense reasoning. It consists of 13k English claims covering 2.7k entities, each labeled as true or false. These claims require reasoning that bridges factual information (e.g., “Harry Potter is a wizard”) with unstated commonsense inferences (e.g., “If someone is good at a skill, they can teach it”). Unlike prior commonsense benchmarks that focus on generic physical or social scenarios, CREAK emphasizes entity-grounded reasoning and assesses whether models can verify claims that depend on both knowledge retrieval and implicit reasoning. Table 10 summarizes the dataset statistics.
Split # Claims Average Length # Unique Entities Vocab Size
Total True False (# tokens)
Train 10,176 5,088 5,088 10.8 2,096 19,006
Dev 1,371 691 680 9.7 531 4,520
Test 1,371 707 664 9.9 538 4,620
Test (Contrast) 500 250 250 10.0 226 1,596
Table 10: Data statistics of CREAK.
A.3 Prompt Details
The following prompt is used to combine the information retrieved from the knowledge graph $G$ (context) and the query logical form $\mathcal{L}(q)$ (query) to form the input to the LLMs discussed in the Section False Premise Detection with Logical Form.
We use the following prompt for logical form conversion:
You are given a question. The task is to: 1) define all the predicates used in the question. 2) parse the question into logic rules based on the defined predicates 3) translate any logical rules implied by the question. 4) convert the question into a logical form using predicate logic. Provide your final answer in the following format: Logical form: Predicate1(entity1, entity2). Keep all expressions concise and consistent. Use standard predicate logic notation.
A.4 Additional Results
A.4.1 Query-level Hallucination Mitigation Analysis
Model Extra Correct FPQ Improved TPQ Change
GPT-4o-mini 427 392 +35
GPT-3.5 50 60 -10
Llama-3.1 139 115 +24
Mistral-7B 94 108 -14
Qwen2.5 139 124 +15
Qwen1.5 94 74 +20
Table 11: Premise-level breakdown of hallucination mitigation improvements. Extra Correct counts queries newly answered correctly compared to direct prompting. FPQ Improved counts false-premise queries corrected after false-premise detection. TPQ Change indicates net changes on true-premise queries after the detection and informing process.
A.4.2 Logical Form Ablation Study
Full w/o Rel. w/o Ent1 w/o Ent2
Logical Form for Retrieval + Original Query for Detection
TPR 0.377 0.044 0.067 0.011
TNR 0.866 0.000 0.067 0.267
FPR 0.133 1.000 0.933 0.733
FNR 0.622 0.956 0.933 0.989
F1 0.540 0.073 0.109 0.020
Acc 0.800 0.038 0.067 0.048
Logical Form for Both Stages
TPR 0.600 0.044 0.067 0.044
TNR 0.866 0.000 0.133 0.267
FPR 0.133 1.000 0.867 0.733
FNR 0.400 0.956 0.933 0.956
F1 0.740 0.073 0.109 0.076
Acc 0.829 0.038 0.076 0.076
Table 12: Logical form ablation results using with RAG.
Full w/o Rel. w/o Ent1 w/o Ent2
Logical Form for Retrieval + Original Query for Detection
TPR 0.822 0.178 0.078 0.111
TNR 0.933 0.133 0.067 0.133
FPR 0.066 0.867 0.933 0.867
FNR 0.177 0.822 0.922 0.889
F1 0.869 0.270 0.124 0.177
Acc 0.838 0.171 0.076 0.114
Logical Form for Both Stages
TPR 0.944 0.167 0.089 0.133
TNR 0.991 0.133 0.133 0.200
FPR 0.009 0.867 0.867 0.800
FNR 0.056 0.833 0.911 0.867
F1 0.971 0.254 0.144 0.211
Acc 0.952 0.162 0.095 0.143
Table 13: Logical form ablation results using G-retriever.
Full w/o Rel. w/o Ent1 w/o Ent2
Logical Form for Retrieval + Original Query for Detection
TPR 0.089 0.056 0.089 0.078
TNR 0.933 0.133 0.200 0.267
FPR 0.067 0.867 0.800 0.733
FNR 0.911 0.944 0.911 0.922
F1 0.162 0.093 0.145 0.130
Acc 0.813 0.067 0.105 0.105
Logical Form for Both Stages
TPR 0.089 0.100 0.100 0.067
TNR 0.933 0.133 0.200 0.267
FPR 0.067 0.867 0.800 0.733
FNR 0.911 0.900 0.900 0.933
F1 0.162 0.163 0.162 0.112
Acc 0.813 0.104 0.114 0.095
Table 14: Logical form ablation results using GraphRAG/ToG.
We examine the contribution of individual components in the logical form, such as entity arguments and relational structure, by analyzing how the removal of specific elements affects false-premise detection across different retrieval backends. We consider variants where the logical form is used for retrieval while the original query is used for false-premise detection, as well as variants where the logical form is applied to both stages. The results are shown in Tab. 12, 13, 14. Across all retrieval backends, removing relational structure or entity arguments from the logical form consistently degrades performance. This suggests that the logical form is most effective when its constituent elements are jointly preserved, and that each component contributes complementary information for identifying unsupported premises.
A.4.3 Comparison with Post-hoc Method
Tab. 18 presents a performance comparison between Contrastive Decoding (Shi et al., 2023), a post-hoc hallucination mitigation method, and other LLMs (Mistral-7B, Qwen1.5, Qwen2.5-7B-Instruct, Llama-3.1-8B-Instruct). Our method achieves improved performance over Contrastive Decoding on all models except Mistral-7B.
A.4.4 False Premise Detection
We additionally evaluate Llama-3.1-8b, as well as GPT-3.5-turbo and G-retriever on the false premise detection task using our method. The results are presented below (Tab. 15 and Tab. 17, Fig. 3). Notably, when original queries are used in either retrieval, false premise detection, or both stages, despite achieving high accuracy (91.11%), G-retriever shows a markedly lower TPR (37.78%) compared to the first configuration. This suggests that relying on original queries alone, or in combination with logical forms in only one stage for detection, can achieve high accuracy due to correctly identifying negatives, it is less effective at capturing false premises, which is the primary focus of our task.
Metric Direct Claim with RAG G-retriever GraphRAG/ToG
Original Query for Both Stages
TPR 0.878 0.811 0.644 0.800
TNR 0.200 0.467 0.467 0.267
FPR 0.800 0.533 0.533 0.733
FNR 0.122 0.189 0.356 0.200
F1 0.873 0.854 0.744 0.832
Accuracy 0.781 0.762 0.619 0.724
TPR 0.878 0.811 0.711 0.800
TNR 0.200 0.400 0.333 0.267
FPR 0.800 0.600 0.667 0.733
FNR 0.122 0.189 0.289 0.200
F1 0.873 0.849 0.780 0.832
Accuracy 0.781 0.752 0.657 0.724
Logical Form for Both Stages
TPR 0.878 0.800 0.811 0.800
TNR 0.200 0.867 0.400 0.267
FPR 0.800 0.133 0.600 0.733
FNR 0.122 0.200 0.189 0.200
F1 0.873 0.878 0.849 0.832
Accuracy 0.781 0.810 0.752 0.724
Table 15: KG-FPQ dataset: comparison of performance metrics across different retrieval methods using original queries and logical forms at different stages using Llama-3.1-8b.
A.4.5 Premise Detection Baseline
We include SAC3 Zhang et al. (2024) as baseline for premise detection for the KG-FPQ dataset. Our proposed approach achieves better performance when considering both F1 score and Accuracy (see Table 2 for comparison).
Method TPR TNR FPR FNR F1 Acc
SAC3 81.1 73.3 26.7 18.9 87.4 80.0
Table 16: Performance of SAC3 on premise detection on KG-FPQ dataset.
Metric G-retriever
Original Query for Both Stages
True Positives (TP%) 37.78
True Negatives (TN%) 100.00
False Positives (FP%) 0.00
False Negatives (FN%) 62.22
F1 Score (%) 54.84
Accuracy (%) 91.11
Logical Form + Original Query
True Positives (TP%) 37.78
True Negatives (TN%) 100.00
False Positives (FP%) 0.00
False Negatives (FN%) 62.22
F1 Score (%) 54.84
Accuracy (%) 91.11
Logical Form for Both Stages
True Positives (TP%) 75.56
True Negatives (TN%) 80.00
False Positives (FP%) 20.00
False Negatives (FN%) 24.44
F1 Score (%) 84.47
Accuracy (%) 79.37
Table 17: False Premise Detection Performance using GPT-3.5-turbo and G-retriever.
Figure 5: Additional comparison of performance metrics across different retrieval methods using logical forms and/or original queries.
Mistral-7B Qwen1.5 Qwen2.5-7B-Instruct Llama-3.1-8B-Instruct
Contrastive Decoding 89.5 76.2 85.7 84.8
Ours 87.6 89.5 92.4 86.7
Table 18: Comparison between contrastive decoding and our method across different LLMs. Note: GPT-3.5 and GPT-4o-mini are not included as logits are not available for contrastive decoding approach.
A.4.6 Case Study